OpenStudy (anonymous):

WHAT IS THE ANSWER TO THIS PROBLEM 5/6-2/3+1/4 DIVIDED BY 3/4+1/3

hero (hero):

a ruby will help you

hero (hero):

without secretly blushing to herself

OpenStudy (anonymous):

OK

OpenStudy (aroub):

=5/6 -2/3+1/4 / 3/4 +1/3 = 5/6*2-2/3*4+1/4*3 / 3/4*3 +1/3*4 = 10/12-8/12+3/12 / 9/12 +4/12 = 5/12 / 13/12 = 5/12*12 / 13/12*12 =5/13

hero (hero):